HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025, Page 3 of 8  Not logged in ELOG logo
ID Date Author Category Subject Text Attachments
  119   Thu Apr 3 11:51:00 2025 CN, IDDAQRSA started for 188Hf ESR tuningStarted both RSAs for the tuning of the ESR
to 188Hf. Both pockets are in parking position
OUTSIDE.
  
  118   Thu Apr 3 10:07:34 2025 CN, ZP, CJG, MCGeneralScreenshots   2025-04-03_10-04-14-352.png2025-04-03_10-04-23-381.png2025-04-03_10-05-06-655.png 
  117   Thu Apr 3 09:53:02 2025 Ruijiu Chen,WolframAnalysisGood injection ratioCurrently, we only use 37% of the injections
in our data analysis. However, 63% of the
injections have a height below 100, and these
 photo_2025-04-03_09-53-26.jpg 
  116   Thu Apr 3 09:43:14 2025 AllAcceleratorStop 194Pb, switch to 208Pb82+The repair of SIS electron cooler power supplies

In the meantime we received a few injections
  
  115   Thu Apr 3 09:34:42 2025 IDGeneralFinished in the ESR as of 9:05We had stopped the RSAs at 8:00 (call for
restart)
Time stamp computer stopped and started at
  
  114   Thu Apr 3 08:02:04 2025 Team PLEIADES (JG, ID, CJG, MC)DetectorsEntering ESR for PLEIADESWe are going to enter the ESR to check on
the PLEIADES pocket detectors and install
a nitrogen cooling.
  
  113   Thu Apr 3 07:59:09 2025 ZP, CN, IDGeneralScreenshotsScreenshots before opening ESR.  2025-04-03_07-58-00-874.png2025-04-03_07-57-08-478.png2025-04-03_07-57-36-781.png 
  112   Thu Apr 3 06:56:31 2025 JEL, ZNRunsScreenshots   2025-04-03_06-55-08-030.png2025-04-03_06-54-58-081.png2025-04-03_06-55-25-792.png 
  111   Thu Apr 3 06:33:08 2025 Zac, JohanGeneralBeam lost at 06:27, returned at 06:34Spike in beam intensity at 06:27 (5x usual
intensity) and then beam is lost. HKR says
it's a problem with the magnet's power supply.
  
  110   Thu Apr 3 06:16:32 2025 JEL, ZNGeneralScreenshotsStable since last entry  2025-04-03_06-15-12-250.png2025-04-03_06-15-00-458.png2025-04-03_06-14-54-153.png 
  109   Thu Apr 3 05:11:37 2025 Zac, JohanGeneralBeam lost at 05:07, returned at 05:14    
  108   Thu Apr 3 03:48:40 2025 JEL, ZNRunsScreenshots
The peaks in the analyzer are slowly shifting
back to the original position. 
 2025-04-03_03-47-28-530.png2025-04-03_03-47-11-337.png2025-04-03_03-47-42-665.png 
  107   Thu Apr 3 02:59:15 2025 JEL, ZNRunsScreenshot, beam received, baem lost, beam back againBeam is back for 10 mins-

The peaks on the analyzer appear shifted
 2025-04-03_02-58-06-109.png2025-04-03_02-57-54-019.png2025-04-03_02-57-31-430.png 
  106   Thu Apr 3 02:52:27 2025 Zac, JohanGeneralBeam is back at 02:52    
  105   Thu Apr 3 02:01:29 2025 Zac, JohanGeneralQuadrupole issuesHKR has called someone to fix the quadrupole,
they think it will take about an hour.
  
  104   Thu Apr 3 01:56:57 2025 Zac, JohanGeneralProblems with one of the quadrupolesSince around 01:45 we get spills but see nothing
on analysers. HKR says there is a problem
with one of the quadrupoles; they are working
  
  103   Thu Apr 3 01:29:49 2025 Zac, JohanGeneralBeam is back at 01:27    
  102   Thu Apr 3 01:19:44 2025 Zac, JohanGeneralBeam fluctuationsBeam gone again between 01:14 and 01:18, then
again at 01:19. HKR says there are problems
with the source.
  
  101   Thu Apr 3 01:14:16 2025 Zac, JohanGeneralBeam is back at 01:12    
  100   Thu Apr 3 00:49:00 2025 Zac, JohanGeneralBeam lost at 00:46    
ELOG V3.1.5-fc6679b